# UniMLVG: Unified Framework for Multi-view Long Video Generation with Comprehensive Control Capabilities for Autonomous Driving

### v2

@@ 4.3 Ablation Studies
We conduct a series of ablation studies to evaluate the distinct contributions of multi-condition, multi-task, multi-dataset and multi-stage training. All the following evaluation metrics are reported on the nuScenes validation set.
@@
$\surd$ 82.26 11.83
$\surd$ $\surd$ 76.52 11.71
$\surd$ $\surd$ $\surd$ 68.80 10.97
@@
Multi-condition. We compare the experimental results for different combinations of the four conditions, as shown in Table 2. We find that incorporating 3D conditions lead to substantial improvement in FVD and FID, as the 3D information helps the model interpret the driving scene. In addition, we also conduct the comparison between implicit (Cam.id.) and explicit (Cam.ray.) camera pose modeling. Compared to Cam.ray., Cam.id. used in  [12] encodes the camera’s intrinsic and extrinsic parameters into a vector, which is handled as global information like textual inputs. The results show that the explicit modeling approach not only improves video continuity by 6.8% but also slightly enhances image quality. This suggests that explicitly incorporating positional relationships between viewpoints into the model enhances its understanding of driving scenes. Additionally, the implicit modeling approach does not improve the model’s performance.
Multi-dataset. Table 3 shows the results of training on the nuScenes [5] dataset with and without pretraining weights from a single-view dataset. The results clearly demonstrate that leveraging a large amount of single-view data significantly enhances the continuity and quality of the videos. This indicates that substantial unlabeled single-view street scene sequential data can effectively enhance the model’s ability to imagine and predict future scenes. Additionally, by integrating two multi-view video datasets, Waymo and Argoverse2, for joint training, we achieved a 6.7% improvement in FVD. This demonstrates the model’s scalability and highlights the critical role of diverse data in enhancing its understanding of driving scenes.
@@ 6 Details on Explicit Perspective Modeling
Explicit perspective modeling aims to inject spatial information into UniMLVG to enhance the coherence of generated videos. Specifically, we utilize a set of the camera’s intrinsic parameters $K\in\mathbb{R}^{B\times T\times V\times 3\times 3}$ and the extrinsic transformations $E\in\mathbb{R}^{B\times T\times V\times 3\times 4}$, which map viewpoints to the forward-facing perspective of the first reference frame. In this way, we can directly obtain the camera’s origin coordinates $Ray\_o=E_{4}$, where $E_{4}\in\mathbb{R}^{B\times T\times V\times 3}$ represent the latest columns of the last dimensions the camera extrinsic matrices. We extend the camera origins to $\mathbb{R}^{B\times T\times V\times 3\times H\times W}$ to match the number of video pixels. We then define a three-dimensional pixel plane (in homogeneous coordinates) $P\in\mathbb{R}^{B\times T\times V\times 3\times H\times W}$ with the same dimensions as the image latent space. Using the scaled camera intrinsic parameters and transformations, the ray directions from the origin to the plane can be computed as follows: $Ray\_d=E_{:3,:3}\times K^{-1}\times P$, $E_{:3,:3}$ refers to the upper-left $3\times 3$ rotation matrix of the extrinsic matrix $E$. After transforming the encoded $Ray\_o$ and $Ray\_d$ through an MLP, the resulting features are added to the image latent before feeding it into the cross-view and temporal modules.
@@ 7 Details on fusion of Cross-view and Temporal Information
The fusion of cross-view and temporal information is essential to this task. We believe that the original text-to-image generation model [10] already possesses strong image generation capabilities, requiring only minor modifications to the image latent to achieve cross-view consistency and temporal coherence. Therefore, we simply use a learnable parameter to perform a weighted summation of the outputs from the cross-view or temporal models with the backbone’s image latent. Specifically, both the cross-view and temporal modules are GPT-2-style self-attention mechanisms [35] to make necessary adjustments to the image latent. The fusion process can be formulated as:
$z_{l}=\text{Sigmoid}(\alpha)\cdot z^{\prime}_{l}+(1-\text{Sigmoid}(\alpha))\cdot\mathcal{F}_{l}(z^{\prime}_{l}),$

### v3

@@ 4.3 Ablation Studies
We conduct a series of ablation studies to evaluate the distinct contributions of multi-condition, multi-task, multi-dataset and multi-stage training. All the following evaluation metrics are reported on the nuScenes validation set. Unless otherwise specified, the ablation experiments are conducted only on Stage 2. Additionally, the weights obtained from the Stage 1 are used as the initial weights for all experiments to enhance training efficiency.
@@
$\surd$ 58.26 8.83
$\surd$ $\surd$ 52.52 8.71
$\surd$ $\surd$ $\surd$ 44.43 6.78
@@
Multi-dataset. Table 3 presents the training results with and without integrating datasets other than the nuScenes [5] dataset. The results clearly demonstrate that leveraging a large amount of single-view data significantly enhances the continuity and quality of the videos. This indicates that substantial unlabeled single-view street scene videos can effectively enhance the model’s ability to imagine and predict future scenes. Additionally, by integrating two multi-view video datasets for joint training, we achieved a 15.4% improvement in FVD. This demonstrates the model’s scalability and highlights the critical role of diverse data in enhancing its understanding of driving scenes.
Multi-condition. Based on the ablation results from multiple datasets, we use the weights trained on the OpenDV-Youtube as initialization for the subsequent experiments. We compare the results for different combinations of the four conditions, as shown in Table 2. We find that incorporating 3D conditions lead to substantial improvement in FVD and FID, as the 3D information helps the model interpret the driving scene. In addition, we also conduct the comparison between implicit (Cam.id.) and explicit (Cam.ray.) camera pose modeling. Cam.id. used in  [12] encodes the camera’s intrinsic and extrinsic parameters into a vector, which is handled as global information like textual inputs. The results show that the explicit modeling approach not only improves video continuity by 12.5% but also enhances image quality by 22.4%. This suggests that explicitly incorporating positional relationships between viewpoints into the model enhances its understanding of driving scenes. Additionally, the implicit modeling approach does not improve the model’s performance.
@@ 6 Details on Explicit Perspective Modeling
Explicit perspective modeling aims to inject spatial information into UniMLVG to enhance the coherence of generated videos. Specifically, we utilize the camera’s intrinsic parameters $K\in\mathbb{R}^{B\times T\times V\times 3\times 3}$ and the extrinsic transformations $E\in\mathbb{R}^{B\times T\times V\times 3\times 4}$ to obtain ray maps that match the size of the images. Importantly, we establish a unified coordinate system with the optical center of the forward-facing camera in the first frame as the origin. In this way, we can obtain the camera’s origin coordinates $Ray\_o=E_{4}$, where $E_{4}\in\mathbb{R}^{B\times T\times V\times 3}$ represent the latest columns of the camera extrinsic matrices. We extend the camera origins to $\mathbb{R}^{B\times T\times V\times 3\times H\times W}$ to match the number of pixels. We then define a three-dimensional pixel index plane (in homogeneous coordinates) $P\in\mathbb{R}^{B\times T\times V\times 3\times H\times W}$ with the same dimensions as the image latent space. Using the scaled camera intrinsic parameters and transformations, the ray directions from the origin to the plane can be computed as follows: $Ray\_d=E_{:3,:3}\times K^{-1}\times P$, $E_{:3,:3}$ refers to the upper-left $3\times 3$ rotation matrix of the extrinsic matrix $E$. After transforming the encoded $Ray\_o$ and $Ray\_d$ through an MLP, the resulting features are added to the image latent before feeding it into the cross-view and temporal modules.
@@ 7 Details on fusion of Cross-view and Temporal Information
The fusion of cross-view and temporal information is essential to this task. We believe that the original text-to-image generation model [10] already possesses strong image generation capabilities, requiring only minor modifications to the image latent to achieve cross-view consistency and temporal coherence. Therefore, we simply use a learnable parameter to perform a weighted summation of the outputs from the cross-view or temporal models with the backbone’s image latent. Specifically, both the cross-view and temporal modules are GPT-2-style self-attention mechanisms [35]. The fusion process can be formulated as:
$z_{l}=\text{Sigmoid}(\alpha)\cdot z^{\prime}_{l}+(1-\text{Sigmoid}(\alpha))\cdot\mathcal{F}_{l}(z^{\prime}_{l}),$ (3)